TMS Meeting Minutes, February 4, 2021

Attendees: Andrew, Breann, Heejoo, John, Matthieu, Trenton, Olga, Yang (Zoom Meeting)

  • No chance to open the telescope last night
  • Olga was able to slew the telescope, create a catalog of a certain position, and slew the telescope there
  • If we are closed again, Olga suggests another dry run
  • Heejoo has a new mic, now Andrew wants one
  • Ch. 8 was not aligned last night (numbers all over the place, signal poor)
  • Andrew using Etalon 3D solution, so we can use both Etalon and our own kinematics; remove faulty ch. 11
  • Short term solution is to get another fiber/collimator set to plug in at ch. 11, as this may solve the problem on ch. 11 which we do not understand
  • Mike visiting LBT next Tuesday to work on SX fiber routing, Heejoo will ask him to look at Ch. 8 and Ch. 11
  • Mike currently has the solid retro, and he is also running a test with our TMS system
  • We may be able to test TMS again on Monday or Tuesday very late at night

Collimators
  • Heejoo reached out to Thorlabs and Edmund Optics engineers regarding changing the collimator-fiber tip distance to suit our needs (still awaiting the response)
  • Since we need to decrease the distance, we may be able to slice a section out of the barrel and reattach
  • Tolerance around 0.3 mm on a 0.8-1.5 mm barrel length decrease, so machining the part could work well

  • Andrew will be reaching out to Etalon soon about the fiber switch and MPO casette
  • Heejoo reached out to Chang-jin Oh about the Leica IR laser tracker
    • Andrew thinks that the Leica may not work with the solid retro, as the wavelengths are incompatible (Leica at 789 nm, retro designed for 1532 nm)
    • API in Germany said that the Radian Laser Tracker uses VIS laser for IF and uses the IR laser for ADM (which is what we are interested in)
    • Still a question as to whether the red laser is required for target locking on ADM mode (some GMT activities indicate this might be the case)
  • Looks like the API laser tracking system is the one to test the solid retro with
GMT Edge Sensor
  • Funny results on the nanometer scale over 900 mm path lengths and over short path lengths, associated with Ciddor correction
  • Also, the attocube interferometer is showing some funny results
  • Andrew carried over some assumptions about temperature having 10 times greater effect on path length than pressure, but now he thinks we may need to worry a little more about pressure
  • At GMT, need to measure micro bars over 900 mm path
  • Breann shared a plot of the Diff M1 length measurements (1 meter) using distance measuring [differential] interferometer, with a Ciddor correction comprised of temperature, humidity, and pressure contributions
  • Retroreflector and collimator are both sitting on Zerodur, so the base is, nominally, not moving at all
  • The Ciddor correction should make this a flat line (but it does not)
  • Humidity has minimal effect on the path length
  • Applying only the temperature contribution, the measured length shifts down ~500 nm
  • Applying only the pressure contribution, the measured length shifts by much more, down ~1500 nm
  • This would indicate that the pressure may have a much larger contribution than previously anticipated
  • By applying the Ciddor correction, we obtain the same result as if we combine the individual effects from temperature, pressure, and humidity contributions
  • Andrew notes that 4 micro bars gives 1 nm path length change, but also 1 mK gives 1 nm path length change, over the same distance (seems incompatible)
  • The pressure contribution is significant
  • Over 9 meter path length, 1 micron path length change corresponds to 0.4 mbar (while we get 4 microns change for a 0 to 100% humidity change, less severe)
  • If a large front were to roll in, could result in a substantial change in pressure, giving a 100 micron change in path length
  • In the case of a significant change of more than a few mbar, need to be sure to apply the Ciddor correction
  • Andrew suggests a historical data mining campain of pulling out temperature, pressure telemetry data from a night when the telescope was working
  • Able to then assess the impact variability in temperature and pressure have on length change
  • Andrew suggests variable scale factor on the temperature and pressure corrections and see if they can be fit to give a flat pseudo-Ciddor correction

  • Temperature, pressure, and humidity data are very different between summer and winter
  • Need to scale the data montly based on the general trend
  • Wondering about the resolution of the weather station barometer

Proposed hysteresis test
  • Settle at Zenith, down to intermediate elevation, take short exposure images while measuring optics position with TMS, down to 30deg elevation and let the telescope settle, then back up to intermediate elevation and repeat short exposures with TMS measurement
  • Andrew proposes settling for 3 minutes instead of 1 minute, based on the hysteresis curves showing >1 min settling
  • John notes that what happens at Zenith happens no matter how long you stay there
  • The most useful data from the hysteresis test was from moving to an intermediate point from both sides
  • Will be nice to have corresponding imagery

Software Updates
  • Finished draft version of the new software with the standard deviation filter
  • Not yet released into official operations
  • This version has significant changes from previous
  • Sending an email to the list with a complete specification of the current behavior of the software, responses to all kinds of data situations (please review)
  • If we accept the specification, we can move forward
  • Outside of the behavior, there are some large changes internally
  • Only tested under isolated simulator
  • The new code conducts a lot of autonomous actions, streamlining the error handling and restarts that used to be passed to John's script
  • Still just a one pass run, so it relies on several files for history
  • Compatible with the current logging system, though it may present some issues
  • Andrew proposes testing the new code with the ASCII channel length data
  • Yang would like to join next time there is a live operation to see if the new code preforms correctly
  • Should probably plan some daytime telescope time for testing the software
  • At present, able to construct all sorts of interesting scenarios in the simulator which we don't see on the telescope
  • Long term plan, hopefully this will be the agreed upon version for some of the decisions
  • We will need some time to test this on the telescope and get to the bottom of channel dropping
  • After this stage, Yang will move to rebuilding the software for usability

  • Olga planning to use LBC tonight for the second half (with wind)
  • Andrew planning to join up to run TMS
  • Also looking forward to getting TMS data when the telescope is in a bad thermal state

-- TrentonBrendel - 04 Feb 2021
Topic revision: r2 - 05 Feb 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback